OpenStudy (anonymous):

x 2 8 --- + --- = ------ x+1 x-1 x^2- 1 Thank You So Much if you help

OpenStudy (agent0smith):

\[\Large \frac{ x }{ x+1}+\frac{ 2 }{ x-1 } = \frac{ 8 }{ x^2-1 }\] You can get common denominators for the first two fractions, like so: \[\Large \frac{ x }{ x+1}*\frac{ x-1 }{ x-1 }+\frac{ 2 }{ x-1 } *\frac{ x+1 }{ x+1 }\]